A child on a slide is traveling 1.02 m/s 2.10 m off the ground. How fast is the child going at the bottom?


1
Expert's answer
2021-04-19T07:21:32-0400

From the law of energy conservation,

"\displaystyle mgh + \frac{mv_0^2}{2} = \frac{mv^2}{2}"

"2gh + v_0^2 = v^2"

"v = \sqrt{2gh+v_0^2} = \sqrt{2 \cdot 9.81 \cdot2.1+1.02^2}=6.5\; m/s."


Answer: 6.5 m/s